OpenStudy (anonymous):

express in simplest form: x^3 -16x

OpenStudy (cwrw238):

first take x out to give x(x^2 - 16) now you factor what's in the parentheses note its the difference of 2 squares: x^2 - 16 can you factor that?

OpenStudy (cwrw238):

use the identity a^2 - b^2 = ( a + b)(a - b) in your case a = x and b = sqrt16 = 4